A length of lumber 12 feet long is cut into 2 unequal lengths. One piece is 3 times as long as the other.
Find the length of the shorter and longer piece.

Answer by macston(5194)

Let S = the shorter length, and L = the longer length = 3S
S+L=12 feet Substitute for L
S+3S=12 feet
4S=12 feet Divide each side by 4
S=3 feet ANSWER 1: The shorter piece is 3 feet
L=3S=3(3 ft)=9 feet ANSWER 2: The longer piece is 9 feet.